a retiree deposited a total of P15,000 in two different savings account. the interest rate on one accoutnt was 4.5% and the interest rate on the other account was 8.0%. at the end of the year, the interest earned on the first account was P250.00 more than the interest earned on the second account. determine the amount deposited in each account.

Answer by mananth(16949)

let one investment be x
and other y
1 x + 1 y = 15000 .............1

4.50 x -8 y = 250 .............2
Eliminate y
multiply (1)by 8
Multiply (2) by 1
8 x 8 y = 120000
4.50 x -8 y = 250
Add the two equations
12.50 x = 120250
/ 12.50
x = 9620
plug value of x in (1)
1 x + 1 y = 15000
9620 + y = 15000
y = 15000 -9620
y = 5380
y = 5380 1
x= 9620
y= 5380
m.ananth@hotmail.ca